Two resistor R and X connected parallel.The ammeter reads a current of 3A,when R is 5ohms, and 6A when R 2A.Determine the value of X


1
Expert's answer
2021-03-14T19:13:00-0400

Explanations & Calculations



  • As the question describes, the setup should be like the attached sketch.



  • The ammeter reads the total current as it says & that is "the voltage applied to the resistor combination/the total resistance".
  • Then by writing equations to the total current, this sum could be solved

"\qquad\qquad
\begin{aligned}
\small i_{total}&=\small i_x+i_R\\
\small 3A&=\small \frac{E}{X}+\frac{E}{5\Omega}\\
&=\small E\Big(\frac{1}{X}+\frac{1}{5}\Big)\cdots(1)
\end{aligned}" For the first combination of resistors

"\qquad\qquad
\begin{aligned}
\small 6A&=\small \frac{E}{X}+\frac{E}{2\Omega}\\
&=\small E\Big(\frac{1}{X}+\frac{1}{2}\Big)\cdots(2)
\end{aligned}" For the second arrangement

  • By (1)/(2),

"\qquad\qquad
\begin{aligned}
\small \frac{1}{2}&=\small \frac{\large{\frac{1}{X}+\frac{1}{5}}}{\large{\frac{1}{X}+\frac{1}{2}}}\\
\small \frac{1}{X}&=\small \frac{1}{10}\\
\small X &=\small \bold{10\Omega}
\end{aligned}"